. An electron accelerated through a potertial of 10,000 V ftom rest has a de-Broglie wave length 'I'. What should be the accelerEting potential so that the wave length is doubled ? (A) 20,000 v (B) 40,000 v (c) 5,000 v (D) 2500 v

Grade:12

2 Answers

Arun
4 years ago
potential difference , V = 10000 volts
 
charge on electron , q = 1.6 × 10^-19 C
 
mass of electron, m = 9.1 × 10^-31 Kg
 
a/c to De-broglie's wavelength,
 
\lambda=\frac{h}{\sqrt{2qVm}}
 
where h is plank's constant. i.e., h = 6.64 × 10^-34 Js
 
so, wavelength = 6.64 × 10^-34/√(2 × 1.6 × 10^-19 × 10000 × 9.1 × 10^-31)
 
= 6.64 × 10^-34/(5.4 × 10^-23) m
 
= (6.64/5.4) × 10^(-34 + 23) m
 
= 1.2296 × 10^-11 m
 
= 12.296 × 10^-12 m or 12.296 pm
 
hence, De-broglie's wavelength is 12.296 pm
Vikas TU
4 years ago
Dear student 
The above ans is incorrect , 
Please do not follow the solution , 
Feel free to ask questions......
Good Luck